Erica earned a total of $66,750 last year from her two jobs. The amount she earned from her job at the store was $1,150 more than three times the amount she earned from her job at the college. How much did she earn (in dollars) from her job at the college?
Answer by ikleyn(52800)

.
    S + C = 66750      (1)   (for the Store and College)

    S = 3*C + 1150     (2)


You can rewrite these two equations in standard form


    S +   C = 66750    (1) 

    S - 3*C =  1150    (2)


Now subtract equation (1) from equation (2). You will get


        -4*c = 1150 - 66750,   or

         4C = 66750 - 1150

         4C = 65600  ==============>  C = 65600/4 = 16400.


ANSWER.  The earning for the college is  $16400.  The rest  66750 - 16400 = 50350  dollars is the earning for the store.


CHECK.   3*C + 1150 = 3*16400 + 1150 = 50350.    ! Precisely correct !

Solved.